# SWE-Effi: Re-Evaluating Software AI Agent System Effectiveness Under Resource Constraints

### v1

@@ 1 Introduction
Leaderboards for benchmarks like SWE-bench jimenez2024swebench provide a foundation for measuring the progress of AI coding assistants in repository-level SWE (software engineering) tasks, such as resolving issues submitted by developers. Yet, the spotlight on these leaderboards shines almost exclusively on a single metric: the resolve rate. Current evaluation paradigms often operate on the implicit assumption of unlimited resources, but companies and developers live in a resource-constrained reality. Was the solution found in minutes, or did it burn through hours of expensive GPU time? Can a team even afford to run this agent on a daily basis? By ignoring such questions, current benchmarks create a disconnect from the realities of practical AI deployment. While our focus is on software engineering, this is a universal problem: a solution from any AI system should be more than correct—it must also be cost-effective.
@@
This perspective shift is critical for two key frontiers in AI-driven software development. First, a trend emerges where agents consume massive LLM tokens (e.g., by test-time compute) to resolve one particular SWE task and earn marginal gains on leaderboards. This raises a critical question about the law of diminishing returns: Is a 1% improvement in resolve rate worth a 5x increase in cost? We argue that this path may not be the most promising direction for building truly scalable and accessible tools. Second, SWE agent scaffolds are increasingly being used as foundations by research teams to train smarter, self-improving models that specialize in SWE tasks through Reinforcement Learning (RL). In long-horizon RLs (like DeepResearch Agent deepresearchagent and SkyRL cao2025skyrl training), they usually have effectiveness requirements so that the RL process will not run forever. In the SWE task scenario, each RL training step, or "rollout," involves setting up the project environment, LLM reasoning, code execution, and test validation. If a SWE agent scaffold is slow or expensive, it hinders the entire RL process. High latency per rollout becomes a significant barrier to effective RL training, limiting progress across the community. For instance, DeepSWE  deepswe2025 developed a custom scaffold to ensure low-latency performance for RL. The Sky-RL cao2025skyrl project uses OpenHands wang2025openhands but requires extensive optimizations to make training feasible. Similarly, SWE-RL wei2025swerl employs the lightweight Agentless-Mini wei2025swerl yet avoids direct execution feedback in RL, likely due to high execution costs. These challenges arise because RL relies on massive-scale trial and error, directly affected by the scaffold’s performance.
@@
To bridge this gap, we introduce the SWE-Effi, a set of new metrics to re-evaluate AI systems in terms of holistic effectiveness scores. While not a new benchmark, we re-evaluate and analyze AI systems [^1] By “AI system” we refer to a single software system that includes an AI model (LLM) with a software scaffold (e.g., agent) working together to solve a given task. on top of the well-established SWE-bench jimenez2024swebench . To be clear, SWE-Effi is designed to complement existing benchmarks like SWE-bench jimenez2024swebench . SWE-Effi offers a holistic perspective on AI system evaluation and highlights performance under resource constraints, emphasizing effectiveness under the cost budget. Moreover, SWE-Effi serves as a valuable indicator for identifying promising foundations in Reinforcement Learning for SWE tasks.
@@
Our primary contribution is a new evaluation perspective, metrics, and a public leaderboard. We conducted an exploratory study, which analyzed five well-known SWE agent scaffoldings (AutoCodeRover zhang2024autocoderover , SWE-agent yang2024swe , OpenHands wang2025openhands , Agentless xia2024agentless , Agentless-Mini wei2025swerl ) paired with three different LLMs (GPT-4o-mini gpt4omini , Llama-3.3-70B llama , Qwen3-32B Qwen3 ) on a representative subset of SWE issues. While we tracked the final resolve rate, our main focus was on evaluating the overall AI system’s time and resource effectiveness in resolving issues and analyzing the characteristics behind it. The findings we present are not intended to be a final, definitive verdict on these five scaffolds. Rather, they serve as an example of the deeper understanding we can unlock by looking at the same problems from a different perspective.
@@ 2.1 AI System Evaluation on Software Engineering Tasks
With the development of LLMs, researchers have begun to investigate whether LLMs can perform more complicated software engineering tasks beyond direct code generation. Jimenez et al. jimenez2024swebench first introduced SWE-bench to evaluate LLMs’ ability to resolve real GitHub issues, which typically involves understanding the issue, localizing files to be edited, and modifying them. Follow up work SWE-bench-Verified swebench-ver then helped to improve quality and robustness of the benchmark through manual evaluation and selection of high-quality instances by human experts. Baterdinov et al. swe-rebench also extended SWE-bench work by providing 21,000 new real-world tasks in order to expand the limited number of high quality SWE data and ensure better evaluation on SWE tasks. At the same time, Zan et al. zan2025multi and Guo et al. guo2025omnigirl proposed SWE task evaluations with more programming languages as well as including multimodal data such as diagrams and images. Beyond GitHub issue resolution, there have been numerous works aimed at using LLMs for various aspects of software engineering. Mündler et al. mundler2024swt proposed SWT-BENCH to evaluate LLM coding agents’ ability to generate test cases that reproduce issues described in natural language, meanwhile, He et al. swe-perf provided SWE-Perf, a benchmark to evaluate LLM proficiency in improving code performance and optimization. Zhao et al. zhao2024gui introduced GTArena to evaluate multimodal LLM agents for end-to-end automated GUI testing. Zhu et al. zhu2025cve proposed CVE-Bench to challenge LLM agents to exploit vulnerable web applications. However, despite these extensive benchmarks evaluating LLM and agent performance, none of them systematically measure the efficiency of the SWE AI systems, often only focusing on the correctness of the generated solutions.
@@ 2.2 LLM-based methods for GitHub Issue Resolution
In recent years, resolving GitHub issues with LLMs has been an extraordinarily active research topic yang2024swe ; zhang2024autocoderover ; antoniades2024swe ; xia2024agentless ; yang2025enhancing ; tao2024magis ; su2025learn ; wang2025openhands ; gao2025trae . Following recent surveys dong2025survey ; yang2025survey , these methods can be categorized into two types: agentic methods and procedural methods. Agentic methods resolve issues by using LLM to independently plan, make decisions, and use tools to complete a task. In this paradigm, Yang et al.yang2024swe proposed SWE-Agent that equips LLM with basic editing and testing tools in a ReAct yao2023react loop, allowing the agent to reason, observe, and interact with the environment by employing a set of tools. Since the original paper, the authors have also released Mini-SWE-Agent mini-swe-agent that massively simplified the original SWE-Agent, giving the agent only one tool for executing bash commands and simplifying the implementation down to 100 lines of python code. Another agentic scaffold, OpenHands wang2025openhands , augments LLMs with a more extensive toolset, including an IPython server, a web browser, and editing tools. It also implements an interaction mechanism that allows human to be involved in the ReAct loop. In comparison to agentic methods that grant an LLM freedom in planning and deciding the execution flow, procedural methods treat issue resolution as a fixed pipeline that invokes LLMs at scripted checkpoints in the process. For instance, Xia et al. proposed Agentless xia2024agentless that first uses hierarchical localization to identify edited code, then creates several candidate patches and chooses one based on test results. Proposed by Zhang et al., AutoCodeRover zhang2024autocoderover localizes suspicious functions using a set of abstract-syntax-tree-based localization tools, generates multiple patches, and then selects one. Yang et al. introduced KGCompass yang2025enhancing , which constructs a knowledge graph from the codebase, repository issues, and pull requests to aid localization, and subsequently generates and selects patches. While there is no consensus on which approach is better, performance and resource efficiency of a scaffold is highly coupled with the strengths and weaknesses of the underlying LLM.
@@ 3.1.1 Derivation of Normalized Inference Time
A key challenge in evaluating systems that rely on external APIs is that raw wall-clock time is an unreliable metric, as it is heavily influenced by external factors such as network conditions and provider-specific hardware. To establish a standardized measure for inference time, we developed a predictive linear regression model. We chose OpenAI’s GPT-4o-mini gpt4omini API as our performance baseline due to its stable and widely representative performance. We trained this model on a dataset of 515,041 individual raw API call logs collected in our experiment in Section 4, partitioned into a 90% training set and a 10% validation set. The model achieved a strong coefficient of determination (R²) of 0.79 on the validation set, indicating a good fit. The resulting equation is defined as:
@@ 4 Experiment Setup
SWE Scaffold Selection and Configuration: We selected five popular, actively maintained open-source SWE scaffolds from the top of the SWE-bench jimenez2024swebench leaderboard (as of May 2025): AutoCodeRover zhang2024autocoderover , OpenHands wang2025openhands , SWE-Agent yang2024swe , Agentless xia2024agentless , and Agentless-Mini wei2025swerl . AutoCodeRover, Agentless, and Agentless-Mini are procedure-based scaffolds that use LLMs to solve software engineering problems in a structured workflow, whereas OpenHands and SWE-Agent are representative agentic-style scaffolds designed for software engineering tasks. They enable LLMs to autonomously plan, edit code, and execute commands by interacting with external tools (e.g., bash), thereby facilitating end-to-end task completion. We configured those scaffolds based on their official guidelines and adhered to their default iteration or generation limits to provide a baseline for "out-of-the-box" performance, except for the termination criteria for SWE-Agent where we adjusted its API budget to a strict \$1 per issue. This forces the agent to operate under the same kind of financial pressure a real team would face. To ensure our time duration measurements were accurate and unbiased, we explicitly disabled any parallel processing capabilities for the five scaffolds. This allowed us to measure the true, sequential processing time of each scaffold’s core logic, making our time-based comparisons direct and fair. To capture the fine-grained data for our analysis, we augmented each SWE scaffold’s logging and added our own profiling code to collect the precise time cost regarding local CPU computation and the backend LLM inference for each AI system.
@@
  • GPT-4o-mini-2024-07-18  gpt4omini : A proprietary model from OpenAI, designed for high efficiency and speed. It serves as a baseline for performance in cost-sensitive scenarios.
  • Llama-3.3-70B-Instruct  llama : A larger, open-source model from Meta. Its 70B parameter size provides strong, general-purpose capabilities and represents a high-performance option within the open-source ecosystem. Note that this model was quantized to FP8 format.
  • Qwen3-32B  Qwen3 : A mid-sized, 32B parameter open-source model from Alibaba, noted for its reasoning abilities. We selected it specifically for its strong reasoning abilities relative to its size. This allows us to evaluate the performance of a more lightweight but capable alternative, which is an attractive profile for resource-constrained environments.
@@
Benchmark: We evaluated all scaffold-model combinations on a focused subset of 50 issues randomly drawn from the well-respected SWE-bench-Verified swebench-ver dataset. To ensure this subset was a fair representation of the whole, we used stratified sampling, preserving the original distribution of issues across different software projects. This approach ensures our findings are representative while keeping the experiment manageable. We will make our subset publicly available on HuggingFace.
@@ 5 Evaluation of AI System SWE Capability and Efficiency
Table 2: Resource performance metrics comparison of SWE AI systems. Taken as the average result across all issue instances (both resolved and unresolved instances) for SWE-bench-Verified. Inf. - normalized inference time, k - thousand. TIN / OUT - input/output tokens, respectively.
SWE Scaffold Base Model Avg CPU Time (s) Avg Inf. Time (s) Avg TIN (k) Avg TOUT (k) Avg LLM Requests
@@
In this section, we present the main experimental observations for the five selected scaffolds paired with three distinct LLMs described in Section  4. Tables  1 and  2 present the Resource Efficiency (Section  3.2) and Core Performance (Section  3.1) metrics, respectively, which form the basis of our SWE-Effi evaluation. All reported values are aggregated and averaged across all issue instances, including both resolved and unresolved attempts, on our curated subset of the SWE-bench-Verified jimenez2024swebench benchmark. Additionally, we provide our analysis and uncover several crucial trends and trade-offs that are vital for anyone looking to build, deploy, or research performant and efficient AI software engineering systems. Lastly, we provide additional evaluation data in the Appendix.

### v2

@@ 1 Introduction
Leaderboards for benchmarks like SWE-bench [1] provide a foundation for measuring the progress of AI coding assistants in repository-level SWE (software engineering) tasks, such as resolving issues submitted by developers. Yet, the spotlight on these leaderboards shines almost exclusively on a single metric: the resolve rate. Current evaluation paradigms often operate on the implicit assumption of unlimited resources, but companies and developers live in a resource-constrained reality. Was the solution found in minutes, or did it burn through hours of expensive GPU time? Can a team even afford to run this agent on a daily basis? By ignoring such questions, current benchmarks create a disconnect from the realities of practical AI deployment. While our focus is on software engineering, this is a universal problem: a solution from any AI system should be more than correct—it must also be cost-effective.
@@
This perspective shift is critical for two key frontiers in AI-driven software development. First, a trend emerges where agents consume massive LLM tokens (e.g., by test-time compute) to resolve one particular SWE task and earn marginal gains on leaderboards. This raises a critical question about the law of diminishing returns: Is a 1% improvement in resolve rate worth a 5x increase in cost? We argue that this path may not be the most promising direction for building truly scalable and accessible tools. Second, SWE agent scaffolds are increasingly being used as foundations by research teams to train smarter, self-improving models that specialize in SWE tasks through Reinforcement Learning (RL). In long-horizon RLs (like DeepResearch Agent [2] and SkyRL [3] training), they usually have effectiveness requirements so that the RL process will not run forever. In the SWE task scenario, each RL training step, or "rollout," involves setting up the project environment, LLM reasoning, code execution, and test validation. If a SWE agent scaffold is slow or expensive, it hinders the entire RL process. High latency per rollout becomes a significant barrier to effective RL training, limiting progress across the community. For instance, DeepSWE  [4] developed a custom scaffold to ensure low-latency performance for RL. The Sky-RL [3] project uses OpenHands [5] but requires extensive optimizations to make training feasible. Similarly, SWE-RL [6] employs the lightweight Agentless-Mini [6] yet avoids direct execution feedback in RL, likely due to high execution costs. These challenges arise because RL relies on massive-scale trial and error, directly affected by the scaffold’s performance.
@@
To bridge this gap, we introduce the SWE-Effi, a set of new metrics to re-evaluate AI systems in terms of holistic effectiveness scores. While not a new benchmark, we re-evaluate and analyze AI systems [^1] By “AI system” we refer to a single software system that includes an AI model (LLM) with a software scaffold (e.g., agent) working together to solve a given task. on top of the well-established SWE-bench [1]. To be clear, SWE-Effi is designed to complement existing benchmarks like SWE-bench [1]. SWE-Effi offers a holistic perspective on AI system evaluation and highlights performance under resource constraints, emphasizing effectiveness under the cost budget. Moreover, SWE-Effi serves as a valuable indicator for identifying promising foundations in Reinforcement Learning for SWE tasks.
@@
Our primary contribution is a new evaluation perspective, metrics, and a public leaderboard. We conducted an exploratory study, which analyzed five well-known SWE agent scaffoldings (AutoCodeRover [7], SWE-agent [8], OpenHands [5], Agentless [9], Agentless-Mini [6]) paired with three different LLMs (GPT-4o-mini [10], Llama-3.3-70B [11], Qwen3-32B [12]) on a representative subset of SWE issues. While we tracked the final resolve rate, our main focus was on evaluating the overall AI system’s time and resource effectiveness in resolving issues and analyzing the characteristics behind it. The findings we present are not intended to be a final, definitive verdict on these five scaffolds. Rather, they serve as an example of the deeper understanding we can unlock by looking at the same problems from a different perspective.
@@ 2.1 AI System Evaluation on Software Engineering Tasks
With the development of LLMs, researchers have begun to investigate whether LLMs can perform more complicated software engineering tasks beyond direct code generation. Jimenez et al. [1] first introduced SWE-bench to evaluate LLMs’ ability to resolve real GitHub issues, which typically involves understanding the issue, localizing files to be edited, and modifying them. Follow up work SWE-bench-Verified [13] then helped to improve quality and robustness of the benchmark through manual evaluation and selection of high-quality instances by human experts. Baterdinov et al. [14] also extended SWE-bench work by providing 21,000 new real-world tasks in order to expand the limited number of high quality SWE data and ensure better evaluation on SWE tasks. At the same time, Zan et al. [15] and Guo et al. [16] proposed SWE task evaluations with more programming languages as well as including multimodal data such as diagrams and images. Beyond GitHub issue resolution, there have been numerous works aimed at using LLMs for various aspects of software engineering. Mündler et al. [17] proposed SWT-BENCH to evaluate LLM coding agents’ ability to generate test cases that reproduce issues described in natural language, meanwhile, He et al. [18] provided SWE-Perf, a benchmark to evaluate LLM proficiency in improving code performance and optimization. Zhao et al. [19] introduced GTArena to evaluate multimodal LLM agents for end-to-end automated GUI testing. Zhu et al. [20] proposed CVE-Bench to challenge LLM agents to exploit vulnerable web applications. However, despite these extensive benchmarks evaluating LLM and agent performance, none of them systematically measure the efficiency of the SWE AI systems, often only focusing on the correctness of the generated solutions.
@@ 2.2 LLM-based methods for GitHub Issue Resolution
In recent years, resolving GitHub issues with LLMs has been an extraordinarily active research topic [8, 7, 21, 9, 22, 23, 24, 5, 25]. Following recent surveys [26, 27], these methods can be categorized into two types: agentic methods and procedural methods. Agentic methods resolve issues by using LLM to independently plan, make decisions, and use tools to complete a task. In this paradigm, Yang et al.[8] proposed SWE-Agent that equips LLM with basic editing and testing tools in a ReAct [28] loop, allowing the agent to reason, observe, and interact with the environment by employing a set of tools. Since the original paper, the authors have also released Mini-SWE-Agent [29] that massively simplified the original SWE-Agent, giving the agent only one tool for executing bash commands and simplifying the implementation down to 100 lines of python code. Another agentic scaffold, OpenHands [5], augments LLMs with a more extensive toolset, including an IPython server, a web browser, and editing tools. It also implements an interaction mechanism that allows human to be involved in the ReAct loop. In comparison to agentic methods that grant an LLM freedom in planning and deciding the execution flow, procedural methods treat issue resolution as a fixed pipeline that invokes LLMs at scripted checkpoints in the process. For instance, Xia et al. proposed Agentless [9] that first uses hierarchical localization to identify edited code, then creates several candidate patches and chooses one based on test results. Proposed by Zhang et al., AutoCodeRover [7] localizes suspicious functions using a set of abstract-syntax-tree-based localization tools, generates multiple patches, and then selects one. Yang et al. introduced KGCompass [22], which constructs a knowledge graph from the codebase, repository issues, and pull requests to aid localization, and subsequently generates and selects patches. While there is no consensus on which approach is better, performance and resource efficiency of a scaffold is highly coupled with the strengths and weaknesses of the underlying LLM.
@@ 3.1.1 Derivation of Normalized Inference Time
A key challenge in evaluating systems that rely on external APIs is that raw wall-clock time is an unreliable metric, as it is heavily influenced by external factors such as network conditions and provider-specific hardware. To establish a standardized measure for inference time, we developed a predictive linear regression model. We chose OpenAI’s GPT-4o-mini [10] API as our performance baseline due to its stable and widely representative performance. We trained this model on a dataset of 515,041 individual raw API call logs collected in our experiment in Section 4, partitioned into a 90% training set and a 10% validation set. The model achieved a strong coefficient of determination (R²) of 0.79 on the validation set, indicating a good fit. The resulting equation is defined as:
@@ 4 Experiment Setup
SWE Scaffold Selection and Configuration: We selected five popular, actively maintained open-source SWE scaffolds from the top of the SWE-bench [1] leaderboard (as of May 2025): AutoCodeRover [7], OpenHands [5], SWE-Agent [8], Agentless [9], and Agentless-Mini [6]. AutoCodeRover, Agentless, and Agentless-Mini are procedure-based scaffolds that use LLMs to solve software engineering problems in a structured workflow, whereas OpenHands and SWE-Agent are representative agentic-style scaffolds designed for software engineering tasks. They enable LLMs to autonomously plan, edit code, and execute commands by interacting with external tools (e.g., bash), thereby facilitating end-to-end task completion. We configured those scaffolds based on their official guidelines and adhered to their default iteration or generation limits to provide a baseline for "out-of-the-box" performance, except for the termination criteria for SWE-Agent where we adjusted its API budget to a strict \$1 per issue. This forces the agent to operate under the same kind of financial pressure a real team would face. To ensure our time duration measurements were accurate and unbiased, we explicitly disabled any parallel processing capabilities for the five scaffolds. This allowed us to measure the true, sequential processing time of each scaffold’s core logic, making our time-based comparisons direct and fair. To capture the fine-grained data for our analysis, we augmented each SWE scaffold’s logging and added our own profiling code to collect the precise time cost regarding local CPU computation and the backend LLM inference for each AI system.
@@
  • GPT-4o-mini-2024-07-18  [10]: A proprietary model from OpenAI, designed for high efficiency and speed. It serves as a baseline for performance in cost-sensitive scenarios.
  • Llama-3.3-70B-Instruct  [11]: A larger, open-source model from Meta. Its 70B parameter size provides strong, general-purpose capabilities and represents a high-performance option within the open-source ecosystem. Note that this model was quantized to FP8 format.
  • Qwen3-32B  [12]: A mid-sized, 32B parameter open-source model from Alibaba, noted for its reasoning abilities. We selected it specifically for its strong reasoning abilities relative to its size. This allows us to evaluate the performance of a more lightweight but capable alternative, which is an attractive profile for resource-constrained environments.
@@
Benchmark: We evaluated all scaffold-model combinations on a focused subset of 50 issues randomly drawn from the well-respected SWE-bench-Verified [13] dataset. To ensure this subset was a fair representation of the whole, we used stratified sampling, preserving the original distribution of issues across different software projects. This approach ensures our findings are representative while keeping the experiment manageable. We will make our subset publicly available on HuggingFace.
@@ 5 Evaluation of AI System SWE Capability and Efficiency
[…]
@@
In this section, we present the main experimental observations for the five selected scaffolds paired with three distinct LLMs described in Section  4. Tables  1 and  2 present the Resource Efficiency (Section  3.2) and Core Performance (Section  3.1) metrics, respectively, which form the basis of our SWE-Effi evaluation. All reported values are aggregated and averaged across all issue instances, including both resolved and unresolved attempts, on our curated subset of the SWE-bench-Verified [1] benchmark. Additionally, we provide our analysis and uncover several crucial trends and trade-offs that are vital for anyone looking to build, deploy, or research performant and efficient AI software engineering systems. Lastly, we provide additional evaluation data in the Appendix.